How many objects can you insert into a hash table before you have a greater than 50% chance to have a collision between any two objects if n=100? Assume all hash values are equally likely.
Answer #1

Considering Entries = N on which there will be more than 50% chance to have collision

Number of ways that objects will be different = 100*99*98*.....(100-N+1)

Number of possible combinations of objects with respect to N entries = 100N

Likelihood of having all objects different = (100*99*98*.....(100-N+1)) / 100N

Likelihood of having two objects same and causing collision = 1 - (100*99*98*.....(100-N+1)) / 100N

so,

= 1 - (100*99*98*.....(100-N+1)) / 100N = 50% = 0.5

= 0.5 = (100*99*98*.....(100-N+1)) / 100N

= 0.5*100N = 100*99*98*....(100-N+1)

This equation is satisfied for N = 13